# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v10

@@ 5 Experiments
For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to (Kim et al., 2018a) for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
@@ 6 Conclusion
In this paper, we proposed a faster gradient variant, termed quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train a logistic regression model. The quadratic gradient introduced in this work can be directly constructed from the Hessian, effectively merging first-order gradient (descent) methods with the second-order Newton’s method. There is a promising potential for the quadratic gradient to accelerate other gradient methods, and it may serve as an alternative or even a replacement for the traditional line-search method.

### v12

@@ 5 Experiments
For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to (Kim et al., 2018a) for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for evaluation and encryption. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
@@ 6 Conclusion
In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain. The proposed quadratic gradient is constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.